# QAEncoder: Towards Aligned Representation Learning in Question Answering System

### v1

@@ A.6.1. Main theorem
    (7) $\mathbb{P}\left(\left|\mathbf{q}^{\top}\mathbf{d}-\cos(\theta)\right|\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{2\mathbf{d}^{\top}\Sigma\mathbf{d}}\right).$
  2. (2)
@@
    (8) $\mathbb{P}\left(\left|\mathbf{q}^{\top}\mu-1\right|\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{2\mu^{\top}\Sigma\mu}\right).$
  3. (3)
@@
    (9) $\mathbf{q}\top\mu-\mathbf{q}\top\mathbf{d}=\mathbf{q}^{\top}\mu(1-\cos(\theta))>0.$
@@
The theorem provides a robust theoretical foundation for the QAEncoder’s capabilities in computing the similarity between documents and queries. These concentration inequalities in equation 7 and equation 8 show that cosine similarities between $\mathbf{q}$, $\mathbf{d}$, and $\mu$ are concentrated around their expected values. Inequality 9 indicates that the similarity between the $\mathbf{q}$ and $\mu$ is always greater than between $\mathbf{q}$ and $\mathbf{d}$. This confirms that using the mean vector as a projection in QAEncoder better captures the semantic relationship between queries and documents. This theoretical result aligns with the experimental findings 1, further validating QAEncoder’s effectiveness .
A.6.2. Proof of Similarity Bounds
@@ A.6.3. Normality Test
We employ the Chi-Squared Q-Q Plot to assess whether the squared Mahalanobis distances 3 conform to the chi-squared distribution. By leveraging Lemma 4, We compare the observed $D^{2}$ values with the theoretical quantiles of the chi-squared distribution to assess the conformity of the data to a high-dimensional Gaussian model. We conclude that close alignment of the sample points along the 45-degree reference line indicates support for the original hypothesis.
Figure 5. Q-Q plot against chi-squared distribution
Direct applications of high-dimensional normality tests, such as the HZ test, often lead to Type I errors in our case. E.g., testing on 768-dimensional normal samples revealed that HZ test demands high ratios of sample size to dimensionality, and it particularly susceptible to Type I errors when the sample size is not sufficiently large. Hence, we opted to perform uni-variate normality assessments on marginal distribution instead.
@@
A.6.4. Some properties of Gaussian distribution
@@
A.7. Missing Figures and Tables
Table 5. Complete retrieval performance across six classical datasets: NQ, SQuAD, ELI5, HotPotQA, MSMARCO, and TriviaQA. Higher is better, the best one is bolded. Hyperparameters including QAEncoder variants and weight terms $\alpha,\;\beta$ are optimized simultaneously for six classical datasets. - denotes default or null values.
Model Param NQ SQuAD ELI5 HotPotQA MSMARCO TriviaQA
MRR@8 MAP@8 NDCG@8 MRR@8 MAP@8 NDCG@8 MRR@8 MAP@8 NDCG@8 MRR@8 MAP@8 NDCG@8 MRR@8 MAP@8 NDCG@8 MRR@8 MAP@8 NDCG@8
@@ A.7. Missing Figures and Tables
bge-large-en-v1.5 - 87.0 83.2 68.0 76.2 76.2 79.9 57.7 55.8 55.8 93.4 90.1 83.3 75.4 75.4 81.3 78.4 75.0 67.5
QAEhyb, $\alpha=0.15,\;\beta=0.5$ 88.2 84.2 69.1 78.2 78.2 81.8 59.1 56.7 57.3 94.2 91.0 83.5 73.8 73.8 79.9 80.2 77.5 70.3
multilingual-e5-large - 86.0 82.3 66.3 86.2 86.2 88.9 39.0 37.8 37.6 95.3 92.8 85.4 73.6 73.6 80 76.8 73.9 66.1
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 86.1 81.9 66.7 84.9 84.9 87.9 46.4 45.6 43.3 94.3 91.0 83.4 70.1 70.1 77.3 79.3 77.0 68.9
gte-base-en-v1.5 - 86.2 82.5 67.6 68.1 68.1 72.4 54.5 52.5 51.6 93.4 89.7 78.4 76.9 76.9 82.4 77.7 74.8 67.2
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 85.5 81.6 67.4 74.8 74.8 78.7 57.0 54.8 55.1 92.9 88.9 76.9 76.0 76.0 81.4 78.8 76.1 69.2
jina-embeddings-v2-small-en - 82.4 79.3 63.3 69.5 69.5 73.6 54.3 52.4 51.6 92.9 88.3 78.1 65.3 65.3 73.3 74.6 72.7 64.4
QAEhyb, $\alpha=0.15,\;\beta=0.5$ 83.3 79.9 64.1 72.4 72.4 76.2 53.8 52.3 51.6 91.7 87.3 76.6 68.2 68.2 75.7 75.9 73.9 65.5
contriever - 78.8 75.0 60.8 64.8 64.8 69.4 51.3 49.2 49.8 89.0 84.6 74.1 55.4 55.4 63.6 71.2 68.6 61.4
QAEemb, $\alpha=0.45$ 84.0 80.2 65.7 74.9 74.9 78.9 55.7 54.0 54.4 89.8 86.2 75.1 67.1 67.1 74.2 76.4 73.9 67.4
mcontriever - 52.1 48.7 37.6 49.3 49.3 54.9 43.0 41.2 40.6 83.0 78.3 67.9 49.6 49.6 56.7 59.4 56.8 49.7
QAEhyb, $\alpha=0.45,\;\beta=0.75$ 61.4 58.5 45.9 64.7 64.7 69.5 51.0 49.1 48.3 85.2 80.9 69.5 65.1 65.1 71.8 70.4 67.6 60.4
bce-embedding-base-v1 - 74.4 70.0 55.5 77.2 77.2 81.3 47.3 46.1 45.0 83.0 79.2 67.8 70.9 70.9 77.7 67.8 65.2 57.3
QAEemb, $\alpha=0.3$ 76.4 72.3 56.8 77.1 77.1 81.1 50.1 48.4 48.4 84.1 79.7 68.0 71.2 71.2 78 69.4 66.5 58.4
text2vec-base-multilingual - 53.8 51.1 36.6 40.9 40.9 45.6 38.2 37.0 34.7 51.2 49.0 37.8 52.4 52.4 59.8 41.8 40.1 33.3
QAEhyb, $\alpha=0.6,\;\beta=0.5$ 66.6 63.7 47.2 56.2 56.2 60.6 43.4 42.1 40.8 72.0 69.2 52.7 67.8 67.8 74.2 57.2 55.5 47.1
@@
Table 6. Comprehensive retrieval performance on the latest datasets FIGNEWS and CRUD-RAG. Higher is better, the best one is bolded. Hyperparameters including QAEncoder variants and weight terms $\alpha,\;\beta$ are optimized simultaneously for six latest datasets. - denotes default or null values.

### v3

@@ A.3.1 Main theorem
    $\mathbb{P}\left(\left|\mathbf{q}^{\top}\mathbf{d}-\cos(\theta)\right|\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{2\mathbf{d}^{\top}\Sigma\mathbf{d}}\right).$ (5)
  2. 2.
@@
    $\mathbb{P}\left(\left|\mathbf{q}^{\top}\mu-1\right|\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{2\mu^{\top}\Sigma\mu}\right).$ (6)
  3. 3.
@@
    $\mathbf{q}\top\mu-\mathbf{q}\top\mathbf{d}=\mathbf{q}^{\top}\mu(1-\cos(\theta))>0.$ (7)
@@
The theorem provides a robust theoretical foundation for the QAEncoder’s capabilities in computing the similarity between documents and queries. These concentration inequalities in Equation 5 and Equation 6 show that cosine similarities between $\mathbf{q}$, $\mathbf{d}$, and $\mu$ are concentrated around their expected values. Inequality 7 indicates that the similarity between the $\mathbf{q}$ and $\mu$ is always greater than between $\mathbf{q}$ and $\mathbf{d}$. This confirms that using the mean vector as a projection in QAEncoder better captures the semantic relationship between queries and documents. This theoretical result aligns with the experimental findings in Fig. 1, further validating QAEncoder’s effectiveness.
A.3.2 Proof of Similarity Bounds
@@ A.3.3 Normality Test
We employ the Chi-Squared Q-Q Plot to assess whether the squared Mahalanobis distances 1 conform to the chi-squared distribution. By leveraging Lemma 2, We compare the observed $D^{2}$ values with the theoretical quantiles of the chi-squared distribution to assess the conformity of the data to a high-dimensional Gaussian model. We conclude that close alignment of the sample points along the 45-degree reference line indicates support for the original hypothesis.
Figure 10: Q-Q plot against chi-squared distribution.
Direct applications of high-dimensional normality tests, such as the Henze-Zirkler test, often lead to Type I errors in our case. E.g., testing on 768-dimensional normal samples revealed that Henze-Zirkler test demands high ratios of sample size to dimensionality, and it particularly susceptible to Type I errors when the sample size is not sufficiently large. Hence, we opted to perform uni-variate normality assessments on marginal distribution instead.
@@
A.3.4 Some properties of Gaussian distribution
@@
Appendix B Query Generation Pipeline
For general purposes, we mainly utilized GPT-4o-mini and the zero-shot prompt in Appendix G for query generation on the latest datasets. For extra experiments on robustness, we also adopt Claude-3-5-Haiku-20241022, Qwen-Turbo-2024-11-01 and Deepseek-V3 as the query predictors. The temperature value is 0.95 and the frequency penalty is 0.1, making a balance between diversity and validity. That is, with a higher temperature value, the generated questions can become too divergent to be answered or supported by the documents.
This pipeline prompt is common and effective for automatic question generation (Zhang et al., 2025; Maity and Deroy, 2024). We demonstrate the diversity of generated queries and the robustness of QAEncoder representations.
B.1 The Diversity of Predicted Query
Statistics Results
Zhang et al. (2025) adopted almost the same prompt and generation process. It showed the diversity and quality of generated queries from six dimensions including question types, question length, context coverage, answerability, uncommonness, and answer length. We excerpt the table of question types, showing the diversity of generated queries.
TriviaQA HotpotQA Llama GPT
T1 Identity/Attribution 34.2 39.7 7.5 15.7
T2 General Knowledge 34.5 15.0 7.3 12.3
T3 Location 12.2 14.3 3.1 4.3
T4 Classification/Categorization 4.3 2.7 2.1 1.5
T5 Specific Fact/Figure 10.5 9.5 18.7 24.1
T6 Comparison/Selection 0.1 6.7 1.0 0.6
T7 Verification/Affirmation 0.1 6.5 0.1 0.2
T8 Descriptive/Characterization 3.0 1.5 43.8 28.7
T9 Event/Outcome 0.2 0.8 14.7 10.1
T10 Sequential/Ordering/Causation 0.9 3.2 1.7 2.4
Others 0.0 0.0 0.0 0.1
Table 6: Percentage of different question types across different datasets
Case Study
To provide a general feeling of diversity, we show the predicted queries by GPT and Claude. For two constructed news benchmarks, the test query is sampled without replacement from GPT’s results; for BEIR, the test queries are from original benchmarks.
Case Study of Diversity - Document: On July 20, 1969, Neil Armstrong landed on the moon to accomplish the Apollo 11 mission, achieved by NASAs planning and the Saturn V rocket. - GPT: [’1. What significant event took place on July 20, 1969?’, ’2. Who was the first person to set foot on the moon?’, ’3. Which organization was responsible for planning the Apollo 11 mission?’, ’4. What was the name of the rocket used to accomplish the Apollo 11 mission?’, ’5. What achievement was accomplished by Neil Armstrong during the Apollo 11 mission?’] - Claude: [’1. On what date did Neil Armstrong land on the moon?’, ’2. What was the name of the mission during which Neil Armstrong landed on the moon?’, ’3. Which organization planned the moon landing?’, ’4. What type of rocket was used to achieve the moon landing?’, ’5. Who was the astronaut who landed on the moon during the Apollo 11 mission?’]
Appendix C Cost-Effectiveness
Briefly, QAEncoder is cost-effective and efficient.
  • Firstly, the cost mainly comes from the query generation process during indexing phase. As shown in Fig. 11, 10 queries with an average length of 6-15 tokens are typically sufficient for ordinary documents.
  • Secondly, GPT-4o-mini is currently the cheapest generative model by OpenAI, priced at \$0.075 per million input tokens and \$0.300 per million output tokens. The most recent Qwen2.5-Turbo by Alibaba is even 3.6 times cheaper than GPT-4o-mini, which outperforms GPT-4 on long-context and mirrors GPT-4o-mini on short-sequence (Qwen Development Team, 2024b). Hence, about 0.1 million documents can be processed within 1 dollar via API call ($\frac{1\mathrm{M}*3.6}{0.3*(10*10)}\approx 0.1\mathrm{M}$). The prices will continually decrease as AI develops.
  • Thirdly, the initial query-centric work for sparse retrievers, DocT5Query, demonstrates T5-base with 0.2B parameters is sufficient for query generation, while no improvement are gained with larger models (Cheriton, 2019). We also confirmed Qwen2.5-0.5B-Instruct as a good query generator in our business implementation. We choose GPT-4o-mini just given its out-of-the-box and comprehensively multilingual support for academic research.
  • Finally, in our data flow settings (batch size=1), 2-3 documents can be processed per second with Qwen2.5-0.5B-Instruct + vLLM + BF16 on a single NVIDIA A100 80GB Qwen Development Team (2024a). For batch processing with higher GPU utility, DocT5Query reports sampling 5 queries per document for 8.8M MSMARCO documents requires approximately 40 hours on a single Google TPU v3, costing only \$96 USD (40 hours × \$2.40 USD/hour) in 2019 (Cheriton, 2019).
Appendix D Dataset Details
  • The BEIR benchmark is a meticulously curated collection of 19 datasets, designed to comprehensively evaluate the generalization capabilities of information retrieval (IR) models across a wide range of heterogeneous tasks. Among the 19 datasets, 15 are publicly available and are used for evaluation in our experiments. The 15 datasets selected from BEIR encompass diverse domains: MSMARCO (Nguyen et al., 2016), TREC-COVID(TRECC.) (Voorhees et al., 2021), NFCorpus (Boteva et al., 2016), Natural Questions (NQ) (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), FiQA18 (Maia et al., 2018), ArguAna (Wachsmuth et al., 2018), Touche20 (Bondarenko et al., 2022), CQADupStack(CQADup.) (Hoogeveen et al., 2015), Quora, DBPedia (Hasibi et al., 2017), SciDocs (Cohan et al., 2020), Fever (Thorne et al., 2018), Climate-Fever(ClimateFe.) (Diggelmann et al., 2020), and SciFact (Wadden et al., 2020).
  • CRUD-RAG is a benchmark specifically designed for evaluating RAG systems, incorporating the latest high-quality news data that were not included in the training phase of the language models. It comprises more than 80K news articles sourced from prominent Chinese news websites, all published after July 2023. From the set of queries generated by GPT-4o-mini for each document, we randomly sample one to serve as the test query. The original news is designated as the evidence documents for recall evaluation, ensuring queries are associated with exactly relevant documents. [^2] LlamaIndex adopts the same common practice and provides templated workflow, i.e. generating queries for recall and rerank evaluation. See the website.
Appendix E Metric Details
  • Mean Reciprocal Rank (MRR): Mean Reciprocal Rank: is a statistic measure used to evaluate the effectiveness of a retrieval system by calculating the reciprocal of the rank at which the first relevant result appears. The mathematical formulation is:
    $\text{MRR}=\frac{1}{|Q|}\sum_{i=1}^{|Q|}\frac{1}{\text{rank}_{i}}$
  • Normalized Discounted Cumulative Gain (NDCG): Normalized Discounted Cumulative Gain is a measure of ranking quality that takes into account the positions of the relevant documents. It is based on the concept of discounting the relevance of documents based on their position in the result list. The mathematical formulation is:
    $\text{NDCG}=\frac{DCG_{p}}{IDCG_{p}},$
    The Discounted Cumulative Gain at position $p$ is given by:
    $DCG_{p}=\sum_{i=1}^{p}\frac{2^{\text{rel}_{i}}-1}{\log_{2}(i+1)},$
    Where: - $\text{rel}_{i}$ is the relevance score of the document at rank $i$.
Appendix F Baseline Details
  • BGE models (Xiao et al., 2024)by BAAI, Jina models (Günther et al., 2023) by Jina AI, E5 models (Wang et al., 2023a) by Microsoft, GTE models (Zhang et al., 2024b) by Alibaba-NLP are the most advanced embedding models, featuring multilingual understanding and task-specific instruction tuning capabilities. We choose both vanilla encoders and QA-specific instruction-tuned encoders for test.
  • DPR models (Karpukhin et al., 2020) by Facebook adopt a bi-encoder architecture. DPR models fine-tuned BERT on pairs of questions and passages without additional pretraining, achieving superior performance compared to traditional methods like BM25.
  • RePAQ models (Lewis et al., 2021) are dense retrievers trained on PAQ (Lewis et al., 2021), a large-scale synthetic corpus of question–answer pairs generated from web texts. We evaluate retriever-multi-base-256, a multilingual variant designed for efficient passage retrieval across multiple languages.
  • Quora DistilBERT models (Reimers and Gurevych, 2019), released by the Sentence-Transformers project, are sentence embedding models based on DistilBERT and fine-tuned on the Quora Duplicate Questions dataset (Thakur et al., 2021) for semantic similarity and duplicate question detection. We evaluate both quora-distilbert-base and quora-distilbert-multilingual, the latter being a multilingual variant trained on parallel corpora covering over 50 languages.
  • Training-based approaches mainly include two types: fine-tuning on QA datasets (domain adaptation) and fine-tuning on multi-task instruction datasets. For fine-tuning on QA datasets, we choose mcontriever-msmarco for test, an enhanced variant of the mcontriever model that has been fine-tuned on the MSMARCO. The second category involves fine-tuning models on multi-task instruction datasets, where distinct prompt prefixes are appended to the input text, enabling the model to effectively differentiate between various tasks. In this category, we test the multilingual-e5-large-instruct (Wang et al., 2024) developed by Microsoft, which leverages synthetic instruction data (Wang et al., 2023a) for fine-tuning. To unveil the catastrophic forgetting issue of training-based methods, we incorporate GPL Wang et al. (2021), which predicted queries, mines hard negative samples, and distills the re-ranker for unsupervised domain adaptation. The method utilize pseudo-queries generated from the target domain as supervision for contrastive learning. Following the original settings, we train the GPL models for 140,000 steps with a batch size of 32.
  • Document-centric methods instruct LLMs to generate a pseudo-document for each query. The pseudo-document aims to capture relevant information but does not correspond to a real document and may contain inaccuracies and hallucinations. Subsequently,the pseudo-document is encoded, and its embedding is utilized to retrieve similar real documents based on vector similarity. We choose HyDE (Gao et al., 2023), Query2Doc (Wang et al., 2023b), and QA-RAG (Kim and Min, 2024) for test. Hyde generates multiple pseudo-documents and fuses their embeddings by mean pooling for retrieval. Similarly, Query2Doc concatenates the query with the generated pseudo-document and performs dense retrieval using the embedding of the combined text. QA-RAG (Kim and Min, 2024) enhances retrieval through a two-way mechanism that utilizes both user query and pseudo-documents for respective retrieval.
Appendix G Instruction Templates
We employ specialized prompts to instruct GPT-4o-mini as the question and pseudo-document generator respectively. Only the English prompts are presented due to LaTeX compilation issues with non-English languages. For BEIR benchmarks, declarative words can serve as valid user queries. We adopt the few-shot in-context learning, which enables the model to generate declarative queries that not only resemble natural inputs but also serve a role analogous to document fingerprints.
Appendix H Hyperparameter Selection
As shown in the main body, $\text{QAE}_{\text{emb}}$ maintains competitive performance with single hyperparameter. Hence, $\text{QAE}_{\text{emb}}$ is recommended for accelerating HP search.
  • Firstly, we believe that the optimal hyperparameters are primarily influenced by the inherent characteristics of the embedding model, i.e. the geometric property of embedding space. Therefore, a one-turn search should be sufficient for a given embedding model. That’s why we optimize hyperparameters simultaneously across multiple datasets.
  • Secondly, the one-turn search can also be accelerated under our framework. Indeed, as Fig. 4 shows, the performance of $\text{QAE}_{\text{emb}}$ empirically follows a consistent trend across various models and datasets: it initially rises and then falls as $\alpha$ increases, peaking between 0.3 and 0.6. This unimodal phenomenon enables ternary search with logarithmic trails rather than brute-force search.
  • Finally, the property of datasets also slightly influences the optimal hyperparameters. Specifically, the optimal $\alpha$ for classical datasets is marginally lower than that for latest datasets (refer to Tables 7 and 8 for details). Therefore, selecting the optimal $\alpha$ based on classical datasets represents a cautious and robust strategy, ensuring consistent improvement across both classical and latest datasets.
Appendix I Challenges of Existing Methods
I.1 Challenges of QAEnaive
  • C1.
    Expanded Index Size. Storing all QA pairs significantly increases the index size, leading to a substantial expansion in storage requirements, especially problematic for large-scale corpora.
  • C2.
  • C3.
    Limited Query Handling. Although storing QA pairs individually can address predicted queries, this approach lacks robustness when confronted with the wide-ranging and diverse nature of potential queries (Alting von Geusau and Bloem, 2020).
I.2 Challenges of Document-centric Methods
Document-centric methods such as HyDE and Query2Doc suffer from not only computation overhead at inference time but also hallucination of pseudo-document generation. Specifically, for user queries, these methods generate pseudo-documents as retrieval queries. However, pseudo-documents suffer from hallucination, especially with rapidly updated knowledge. A concrete example hallucination can be:
Pseudo-Document Generator Prompt - User Query: Who won the 2024 Abel Prize? - Pseudo Document (Hallucinatory Retrieval Query): The 2024 Abel Prize was awarded to Abel for proving that it is impossible to solve the general equation of the fifth degree using radicals. - Retrieved Document: Abels most famous single result is the first complete proof demonstrating the impossibility of solving the general quintic equation in radicals. This question was one of the outstanding open problems of his day, and had been unresolved for over 250 years.[2] He was also an innovator in the field of elliptic functions and the discoverer of Abelian functions. - LLM Answer: Abel - Groundtruth: Michel Talagrand
Appendix J More Figures and Tables
Figure 11: The convergence speed for Monte Carlo estimation of QAEbase, $n$ denotes the number of prediction queries. For documents with a length greater than 150 words from MSMARCO datasets, generating 10 queries exhibits a similarity score of 0.96 compared to generating 80 queries. Besides, the document fingerprint strategies introduces the hyperparameter $\alpha$ to QAEbase, further reducing the variance.
Figure 12: Ablation on $\beta$ hyperparameter for QAEtxt on FIGNEWS(English) dataset.
Figure 13: Ablation on $\alpha$ and $\beta$ hyperparameters for QAEhyb on FIGNEWS(English) dataset.
Model Method ArguAna ClimateF. CQADups. DBPedia FEVER FiQA18 HotpotQA MSMARCO NFCorpus NQ Quora SciDocs SciFact Touche20 TRECC.
@@ I.2 Challenges of Document-centric Methods
dpr - 17.5 14.8 15.3 26.3 56.2 11.2 39.1 17.7 18.9 47.4 24.8 7.7 31.8 13.1 33.2
QAEemb, $\alpha=0.45$ 29.6 27.6 22.3 32.4 70.9 20.3 43.4 25.4 23.9 52.7 28.9 13.5 40.4 22.9 48.0
contriever - 37.9 15.5 28.4 29.2 68.2 24.5 48.1 20.6 31.7 25.4 83.5 14.9 64.9 19.3 27.4
QAEemb, $\alpha=0.45$ 47.1 20.2 33.5 33.7 73.1 30.4 50.3 26.1 37.9 28.1 86.2 17.3 70.2 25.2 45.2
contriever-msmarco - 44.6 23.7 34.5 41.3 75.8 32.9 63.8 40.7 32.8 49.8 86.5 16.5 67.7 20.4 59.6
QAEhyb, $\alpha=0.3,\;\beta=0.75$ 53.9 25.5 38.2 44.9 82.0 37.9 65.3 39.9 37.6 52.7 89.9 18.7 73.6 26.3 70.6
bge-large-en-v1.5 - 63.5 36.6 42.2 44.1 87.2 45.0 74.1 42.5 38.1 55.0 89.1 22.6 74.6 24.8 74.8
QAEhyb, $\alpha=0.15,\;\beta=0.5$ 68.8 38.2 45.6 44.8 91.5 48.2 73.9 41.2 41.7 56.4 90.1 25.6 78.9 28.1 78.2
multilingual-e5-large - 54.4 25.7 39.7 41.3 82.8 43.8 71.2 43.7 34.0 64.1 88.2 17.5 70.4 23.1 71.2
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 61.1 28.4 44.3 42.9 82.1 45.1 69.9 43.0 35.5 65.3 88.0 20.8 73.9 26.3 75.1
e5-large-v2 - 46.4 22.2 37.9 44.0 82.8 41.1 73.1 43.5 37.1 63.4 86.8 20.5 72.2 20.7 66.5
QAEhyb, $\alpha=0.3,\;\beta=1.0$ 55.1 25.3 41.2 45.5 86.5 43.4 73.9 42.8 39.8 64.8 89.5 23.3 75.3 23.8 74.2
gte-base-en-v1.5 - 63.5 40.4 39.5 39.9 94.8 48.7 67.8 42.6 35.9 53.0 88.4 21.9 76.8 25.2 73.1
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 68.2 43.2 43.7 39.6 94.2 51.6 66.7 41.9 38.7 52.3 88.2 24.8 80.3 29.3 77.5
jina-embeddings-v2-small-en - 46.7 24.0 38.0 32.7 68.0 33.4 56.5 37.3 30.4 51.6 87.2 18.6 63.9 23.5 65.2
QAEhyb, $\alpha=0.15,\;\beta=0.5$ 55.3 27.8 42.3 35.6 74.3 36.5 55.7 39.5 32.7 51.1 89.9 21.2 67.2 27.2 73.6
@@
Table 7: Complete retrieval performance across fifteen classical datasets in BEIR (NDCG@10). Higher is better, with the best one is bolded. Hyperparameters including QAEncoder variants and weight terms $\alpha,\;\beta$ are optimized simultaneously. ‘-’ denotes default or null values.